OpenStudy (anonymous):

evaluate [(30+6)-3²]/9*2

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

(36-9)/18 27/18 1.5 Got It?

thomaster (thomaster):

@Game&twerk You know how to write equations in latex?

thomaster (thomaster):

Like this: \(\huge\frac{30+6-3^2}{9*2}\) \(\huge\frac{36-3^2}{9*2}~\to~\huge\frac{36-9}{9*2}~\to~\huge\frac{27}{9*2}\) \(\huge\frac{\boxed{\LARGE27}}{\boxed{\LARGE 9}*2} ~\to~\Huge\frac{3}{2}\)